# Large Model Empowered Metaverse: State-of-the-Art, Challenges and Opportunities

### v1

@@ IV-D Performance Evaluation
We conduct simulations to validate the effectiveness of the proposed Metaverse rendering optimization system. We implement the diffusion model with the Attention U-Net architecture [13], configured with 512 dimensions and 8 attention heads. For the training phase, we employ the Adam optimizer with an initial learning rate of 0.0001 and a batch size of 256, training for 20 epochs with early stopping to prevent overfitting. The diffusion steps are set to 700, with a diffusion variance schedule coefficient linearly increasing from 0.0001 to 0.04. For the validation phase, our prototype is implemented on a cloud server running Ubuntu 20.04 OS, equipped with a 2.60 GHz Intel Xeon CPU, 256 GB of memory, and dual NVIDIA RTX 3090 GPUs. We apply two mainstream optimization strategies, i.e., Markov Decision Processes (MDP) and Random Optimization (RO) as benchmarks for performance comparison.
@@
We first evaluate the Metaverse rendering optimization performance in different schemes. As depicted in Fig. 4, the proposed scheme outperforms the benchmarks in accuracy, recall, and F1-Score. Specifically, our proposed scheme reaches the following improvements over the best-performing benchmark: a 13.1% increase in optimization accuracy, a 71.11% boost in optimization recall, and a 47.54% raise in optimization F1-Score. These improvements in optimization performance demonstrate the effectiveness of our proposed scheme in predicting users’ intentions and preferences in Metaverse, adapting to dynamic and personalized rendering requirements, thereby enhancing overall QoS and QoE for Metaverse users.
Next, we evaluate the rendering time of 20 Metaverse scenes on dual NVIDIA RTX 3090 GPUs in different schemes. As observed in Fig. 5, under non-optimization scheme, the Metaverse rendering overhead incurs an average of 39.54 second in 20 different Metaverse scenes. By contrast, the proposed scheme reduces the rendering overhead to an average of 18.14 second, which achieves the best rendering efficiency among all benchmarks, reducing rendering time by approximately 18.44% compared to MDP, 38.38% compared to RO, and 54.12% compared to non-optimization.

### v2

@@ IV-D Performance Evaluation
We conduct simulations to validate the effectiveness of the proposed Metaverse rendering optimization system. We implement the diffusion model with the Attention U-Net architecture [13], configured with 512 dimensions and 8 attention heads. For the training phase, we employ the Adam optimizer with an initial learning rate of 0.0001 and a batch size of 256, training for 20 epochs with early stopping to prevent overfitting. The diffusion steps are set to 700, with a diffusion variance schedule coefficient linearly increasing from 0.0001 to 0.04. For the validation phase, our prototype is implemented on a cloud server running Ubuntu 20.04 OS, equipped with a 2.60 GHz Intel Xeon CPU, 256 GB of memory, and dual NVIDIA RTX 3090 GPUs. We apply two mainstream optimization strategies, i.e., Markov Decision Processes (MDP) and Random Optimization (RO) as benchmarks for performance comparison. All the validation benchmarks are using the optimal parameters according to their respective literature. For the MDP baseline, the rendering policy is optimized using value iteration with a discount factor of 0.95, and for the RO baseline, 21 rendering parameter combinations are randomly sampled per scene, with the best-performing configuration selected.
@@
We first evaluate the Metaverse rendering optimization performance in different schemes. As depicted in Fig. 4, the proposed scheme outperforms the benchmarks in accuracy, recall, and F1-Score. Specifically, our proposed scheme reaches the following improvements over the best-performing benchmark: a 13.1% increase in optimization accuracy, a 71.11% boost in optimization recall, and a 47.54% raise in optimization F1-Score. These improvements in optimization performance demonstrate the effectiveness of our proposed scheme in predicting users' intentions and preferences in Metaverse, adapting to dynamic and personalized rendering requirements, thereby enhancing overall QoS and QoE for Metaverse users.
Next, we evaluate the rendering time of 20 one-minute Metaverse scenes, each comprising 3,600 frames at 60 FPS, on dual NVIDIA RTX 3090 GPUs in different schemes. As observed in Fig. 5, under non-optimization scheme, the Metaverse rendering overhead incurs an average of 39.54 second in 20 different Metaverse scenes. By contrast, the proposed scheme reduces the rendering overhead to an average of 18.14 second, which achieves the best rendering efficiency among all benchmarks, reducing rendering time by approximately 18.44% compared to MDP, 38.38% compared to RO, and 54.12% compared to non-optimization.
@@ VI Conclusion
Acknowledgement
This work was supported in part by NSFC under Grant 62302387, Grant U22A2029, Grant U24A20237, Grant 62302288, Grant 62402379, Grant 62371280.